Suppose that the width of a rectangle is 5 inches shorter than the length and that the perimeter of the rectangle is 50 inches. Set up an equation for the perimeter involving only L, the length in the rectangle.
Also solve this equation algebraically to find the length of the rectangle. Find the width as well.

Answer by checkley77(12844)

W=L-5
P=2L+2W
50=2L+2(L-5)
50=2L+2L-10
4L=50+10
4L=60
L=60/4
L=15 ANS.
W=15-5=10 ANS.
PROOF:
50=2*15+2+10
50=30+20
50=50